OpenStudy (anonymous):

find the linearization L(x) of the function at a. f(x)=sin x, a=pi/6

OpenStudy (thejax):

This is normally the same as finding the tangent line at pi/6 y = f(a) + f'(a)*(x-a) y = 1/2 + sqrt3/2(x-pi/6) y = sqrt3/2 x + (6-pi*sqrt3)/12

OpenStudy (anonymous):

where does the sqrt3 come from?